OpenStudy (anonymous):

Solve the inequality for x, assuming that a, b, and c are positive constants. (a) a(bx − c) ≥ bc (Choose the correct answer) 1.[0, c/a] 2. (c/b, bc/a) 3. [c/a + c/b, infinity) 4. (bc, a(b-c)/bc] 5. (-infinity, c/b]

OpenStudy (anonymous):

\[a(bx-c)\ge bc\\abx-ac\ge bc\\abx\ge ac+bc\\x\ge \frac{ac}{ab}+\frac{bc}{ab}\\x\ge\frac{c}{b}+\frac{c}{a}\]

OpenStudy (anonymous):

Thank you so much! It makes a lot more sense now